While grocery shopping, you are trying to decide which orange juice bottle would be the best to buy in terms of how much real juice is in the bottle compared to water. Both bottles each contain 1 Litre of solution. In Bottle 1, there is 5% real juice. In Bottle 2, there is 200mL of real juice. Which bottle of juice would be the best to buy if you wanted a solution that had the highest concentration of real juice.
1
Expert's answer
2020-11-11T16:14:10-0500

What is the concentration of a juice in Bottle 2?

"\begin{matrix}
 Solution & 1000\ mL & 100\% \\
 Juice & 200\ ml & x\%
\end{matrix}"

Then


"x\%=\dfrac{200ml}{1000mL}\cdot100\%=20\%"

SInce "20\%>5\%" , then Bottle 2 of juice would be the best to buy if you wanted a solution that had the highest concentration of real juice.


Or

How many millilitres of real juice are there in Bottle 1?

"\begin{matrix}
 Solution & 1000\ mL & 100\% \\
 Juice & y\ ml & 5\%
\end{matrix}"

Then


"y \ mL=\dfrac{5\%}{100\%}\cdot1000\ mL=50\ mL"

SInce "200\ mL>50\ mL" , then Bottle 2 of juice would be the best to buy if you wanted a solution that had the highest concentration of real juice.